OpenStudy (josee):

A certain function f is known to have the properties lim x to -infinity f(x) = 7, lim x to infinity f(x) = 8 . Determine if lim x to 0+ of (3 + 6x)/(5 + f (1/x)) exists, and if it does, compute its value.

OpenStudy (josee):

i have no clue as to how the top is relevant

OpenStudy (anonymous):

because if \[\lim_{x\rightarrow \infty}f(x)=8\] then \[\lim_{x \rightarrow 0^+}f(\frac{1}{x})=8\] as well

OpenStudy (anonymous):

since as \[x\rightarrow 0^+\] \[\frac{1}{x} \rightarrow \infty\]

OpenStudy (josee):

ohhhhhhhhhhhhhh i see i see

OpenStudy (josee):

Thank youu

OpenStudy (anonymous):

yw